KTeV Shift Summary -- DAY 9/19/99 (Sun.)

Earl C. Swallow (earls@elmhurst.edu)
Sun, 19 Sep 1999 15:27:29 -0500 (CDT)


Shift Summary for DAY shift, 09/19/99 (Sun.)

On Shift: Jason LaDue & Earl Swallow/Ashkan + BobT, Valeri

Hours of beam available at all intensities: 3.7 (out of 8.5)
Hours of beam available at nominal intensity: 3.7
Hours of E799 data (special runs not included): 3.6
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time & intensity:

Aborts, Switchyard Rad. Monitor failure, MI rf problems,
Booster water leak.

========================================================================
Runs ended this shift:

14626 -- BEAM_799 -- 1.6M
14627 -- BEAM_799 -- 0.3M
14628 -- BEAM_799 -- zip
14629 -- BEAM_799 -- zip
14630 -- BEAM_799 -- zip
14631 -- LASER_SCAN -- aborted/beam returns
14632 -- BEAM_799 -- 3.5M

Tapes brought to FCC this shift:

UP(A,B,C)004-006, UPS002 (runs through 14625)

========================================================================
Accesses during shift:

None.

========================================================================
Overview of shift:

Initial data taking at 8E12 with 10" Be absorber combined with
updating parameters, etc. Lots of accelerator problems.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
07:00 Shift Begins.
Tape change from OWL shift is under way for
end of Run 14625.

14626 07:10 Begin Run 14626 -- BEAM_799.

14626 08:00 End Run 14626 to update TRD constants.

14627 08:08 Begin Run 14627 -- BEAM_799.

08:09 7 spills lost to aborts.

08:15 TRD alignment constants were successfully updated
by Valeri.

08:25 End Run 14627 to edit/fix some triggers (KTeV Log
XXI, pages 242-243).

14628 08:41 Start Run 14628 with revised trigger definitions,
but Beam OFF -- Switchyard Rad. Monitor failure
09:08 Beam returns, but not really (2 weak spills).

14628 09:37 End Run 14628 to update CsI gain constants.
Still no beam.

14629 09:55 Start Run 14629.

14629 10:01 End Run 14629 to fix online plots.

10:11 Mount UPM001 on KTEV4/tps3d2 for Onsplit.

10:09 MI rf problem keeps beam OFF.

14630 10:20 Start Run 14630 -- still no beam.

14630 12:06 End Run 14630 -- no data.

14631 12:10 Start LASER_SCAN (Run 14631).

14631 12:23 Abort Run 14631 because beam returns.

14632 12:31 Start Run 14632 -- BEAM_799.

90 14:28 Beam dies -- Booster water leak.

15:00 ECS has reviewed Monitoring Histograms --
1.5 pages of notes in Logbook.

14632 15:29 End Run 14632 to add 5x prescale to Trigger 6.

15:30 Shift Ends. Still no beam.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

Monday, DAY : Get NEW HP4Si toner cartridges from stockroom and put the
old ones (at least 4 in control room) into the FNAL recycling system.

--
===============================================================================
Earl C. Swallow                           email:  earls@elmhurst.edu
Theophil W. Mueller Professor              (or)   earls@fnal.gov
   and Chairperson
Department of Physics                     Phone:  (630)-617-3577
Elmhurst College                          FAX:    (630)-617-3735
190 Prospect Avenue
Elmhurst, IL   60126-3296                 http://www.elmhurst.edu/~earls
===============================================================================
"Books are lighthouses erected in the dark sea of time."
                                --  from Gargoyles  (The Scrolls of Merlin)